### v1

Abstract
While Multimodal Large Language Models (MLLMs) have experienced significant advancement on visual understanding and reasoning, their potentials to serve as powerful, flexible, interpretable, and text-driven models for Image Quality Assessment (IQA) remains largely unexplored. In this paper, we conduct a comprehensive and systematic study of prompting MLLMs for IQA. Specifically, we first investigate nine prompting systems for MLLMs as the combinations of three standardized testing procedures in psychophysics (i.e., the single-stimulus, double-stimulus, and multiple-stimulus methods) and three popular prompting strategies in natural language processing (i.e., the standard, in-context, and chain-of-thought prompting). We then present a difficult sample selection procedure, taking into account sample diversity and uncertainty, to further challenge MLLMs equipped with the respective optimal prompting systems. We assess three open-source and one close-source MLLMs on several visual attributes of image quality (e.g., structural and textural distortions, color differences, and geometric transformations) in both full-reference and no-reference scenarios. Experimental results show that only the close-source GPT-4V provides a reasonable account for human perception of image quality, but is weak at discriminating fine-grained quality variations (e.g., color differences) and at comparing visual quality of multiple images, tasks humans can perform effortlessly.
Keywords:
Image quality assessment Multimodal large language models Model benchmarking
Figure 1: Illustration of several visual attributes of image quality in our experiments.
1 Introduction
The evolution of Large Language Models (LLMs) has marked a significant milestone in the field of Artificial Intelligence (AI) [53, 49, 6, 2]. The underlying idea of scaling the model size and training data [20] has rendered LLMs the abilities to perform various Natural Language Processing (NLP) tasks with unprecedented levels of accuracy [2].
In the midst of these developments, a particularly promising offshoot has emerged in the form of Multimodal LLMs (MLLMs). These advanced models have taken the capabilities of LLMs a step further by incorporating visual data alongside text [31, 64, 13, 4, 46, 61]. MLLMs typically integrate visual data via Vision Transformers (ViTs) [14] for feature extraction, attention mechanisms [50] for visual-textual relationship modeling, and fusion modules [3, 26, 17] to merge the two modalities. These techniques enable MLLMs to process both text and image data holistically, extending the application scenario of LLMs.
Apart from their high-level visual understanding and reasoning capabilities [31, 61], MLLMs also open up substantial opportunities for Image Quality Assessment (IQA) [52]. As a fundamental vision task, IQA aims to devise computational models to predict image quality as perceived by the Human Visual System (HVS). Ideally MLLMs shall benefit IQA in the following aspects.
  • Improved accuracy. MLLMs are commonly built upon strong visual encoders [42, 32], which are exposed to massive images of various vision tasks closely related to (and including) IQA. This allows MLLMs to cross-validate visual information from different vision tasks for improved IQA, which is in a similar spirit to multitask learning for IQA [71, 35].
  • Improved robustness. Being sequential in nature, MLLMs do not rely on alignment of images for quality assessment, making them easily robust to perturbations that are imperceptible to the human eye, such as mild geometrical transformations [34] and texture substitution [11].
  • Flexibility. Text as one modality of input, MLLMs enable a wide range of text-conditioned IQA, for example, IQA for semantically meaningful local regions [66], fine-grained visual attributes (e.g., color appearance and well-exposedness) [9], and various viewing conditions and display systems [10].
  • Interpretability. MLLMs generate descriptive text rather than merely providing a numerical score [67]. This text output allows for more detailed, contextually rich, and human-like quality assessment, making MLLMs valuable for IQA model-in-the-loop image enhancement [29].
Previous work [55, 56, 67] focuses primarily on establishing IQA datasets with human quality descriptions to benchmark MLLMs in terms of quality question answering, rating, and reasoning. Appealing at first glance, these studies may suffer from several limitations. First, how to properly instruct human subjects to supply detailed and balanced descriptions of image quality and other relevant visual information is highly nontrivial. Second, image quality is a perceptual quantity with subjective biases. How to screen outlier descriptions and aggregate valid but relatively inconsistent descriptions are overlooked. Third, comparing MLLM outputs to reference descriptions is complex, and remains an intriguing and challenging problem in language modeling [28, 1]. After all, the field of IQA has a rich history, and numerous established human-rated image quality datasets are readily accessible for evaluating this perceptual aspect of MLLMs. Closest to ours, Zhu et al. [73] tested the pairwise quality comparison capability of MLLMs in a constrained context.
In this work, we further pursue this direction, and conduct a comprehensive and systematic study of prompting MLLMs for IQA, as opposed to optimizing individual text prompts. We first explore nine prompting systems for MLLMs, combining standardized testing procedures in psychophysics (i.e., the single-stimulus, double-stimulus, and multiple-stimulus methods) with popular prompting strategies in NLP (i.e., the standard, in-context, and chain-of-thought prompting). Interestingly, we discover that the optimal prompting system varies between open-source [31, 64, 13] and closed-source MLLMs. To further challenge MLLMs, we propose a computational procedure to select difficult samples using top-performing IQA expert models as proxies [60, 71, 11, 69, 57, 8], while taking into account sample diversity and uncertainty. Under both Full-Reference (FR) and No-Reference (NR) settings, we experiment with three open-source and one close-source MLLMs across several visual attributes of image quality, including structural and textural distortions, color differences, and geometric transformations (see Fig. 1). Experimental results show that even the most advanced MLLM (i.e., the proprietary GPT-4V) still has room for improvement in modeling the incredibly sophisticated adaptable HVS for IQA.
2 Related Work
In this section, we present a summary of expert models and MLLMs for IQA.
2.1 Expert Models for IQA
IQA can be roughly divided into two categories: FR- and NR-IQA. FR-IQA models are preferred in situations where the unimpaired reference image is available, such as in many low-level vision tasks. Representative design philosophies for FR-IQA range from measuring error visibility (e.g., the mean squared error (MSE)), structural similarity (e.g., the SSIM index [52]), and mutual information (e.g., the VIF measure [43]) to (deep) learning-based methods (e.g., the LPIPS [69] and DISTS metrics [11]) and fusion-based approaches (e.g., VMAF [48]). Almost all FR-IQA models depend on the proper alignment of the reference and test images to execute co-located comparison at the pixel, patch, or feature level. Consequently, these models may struggle to capture the robustness of the HVS to mild geometric transformations and texture resampling. Additionally, these models frequently disregard or give a superficial treatment of color information, leading to a poor account for perceptual color differences [51, 9].
NR-IQA models [37, 70, 60, 21, 71, 59, 63] are the most challenging yet widely applicable as they are designed to evaluate image quality without any reference. Regardless of whether they are knowledge-driven [37] or data-driven [70, 71], NR-IQA models rely heavily on the training data, which often demonstrate weak generalization when faced with a broader type of images and distortions.
Apart from accuracy and generalization, expert IQA models also fall short in terms of flexibility and interpretability. This is primarily because they reduce the evaluation of image quality to a numerical score, failing to leverage other formats of inputs, and output other detailed and perceptually relevant information.
2.2 MLLMs for IQA
Previous work has been centered on benchmarking and refining MLLMs for IQA. The pioneering datasets, DepictQA [67] and Q-Bench [55], have initiated the assessment of MLLMs in quality rating and reasoning under the FR and NR conditions, respectively. DepictQA employs a double-stimulus method, prompting MLLMs to compare a pair of images against a standard reference. In contrast, Q-Bench takes a single-stimulus approach, assigning a numerical score based on the classification of image quality as “poor” or “good”. Notably, these investigations reveal that, except for the proprietary GPT-4V [61], open-source MLLMs like LLaVA [31], Kosmos-2 [41], and MiniGPT-4 [72] show limited success in replicating human perception of image quality. Efforts to enhance open-source MLLMs through instruction tuning have been made [56], focusing on the prediction of subsequent tokens. Yet, such models often revert to generating template-like quality descriptions, lacking the level of flexibility we are looking for. In this work, we take a step back and conduct a comprehensive and systematic exploration of prompting techniques for MLLMs on existing human-rated IQA datasets.
Figure 2: Three standardized psychophysical testing procedures for IQA. (a) Single-stimulus method. (b) Double-stimulus method. (c) Multiple-stimulus method.
3 Systematic Prompting MLLMs for IQA
In this section, we introduce several prompting systems for MLLMs in the context of IQA, followed by describing the difficult sample selection procedure.
3.1 Systematic Prompting from Psychophysics
Although MLLMs and the HVS (or more generally the human brain) operate in different domains, they have fascinating similarities, especially in representational hierarchy [5], visual-textual integration, contextual reasoning [22], and adaptation [12]. Thus, it is natural to employ standardized psychophysical testing procedures as the systematic prompting methods for IQA.
Single-stimulus Method. As depicted in Figure 2 (a), given a test image $x\in\mathcal{X}$, the FR single-stimulus method [^1] This is also known as the double-stimulus impairment rating by treating the reference image as a second visual stimulus. suggests MLLMs to output a quality score, $q(x)\in[0,100]$, based on its perceptual difference to the corresponding reference image $y\in\mathcal{Y}$. Here, a larger number indicates higher quality. In contrast, the NR single-stimulus method derives the quality score solely from the perceptually important visual attributes of the test image $x$. Generally, the single-stimulus method is scalable and straightforward to implement for MLLMs, but may not capture relative quality accurately and may exhibit context dependency (e.g., the presentation order).
Double-stimulus Method. Also known as two-alternative forced choice (2AFC) and paired comparison, the double-stimulus method first samples a pair of images either uniformly or actively [62] from a set of images $\mathcal{X}=\{x_{i}\}^{M}_{i=1}$, where $M$ is the number of images. In our implementation, MLLMs are not forced to select between the two alternatives with higher quality. Instead, they have a third option: the quality of the two images is comparable. Figure 2 (b) illustrates the FR and NR double-stimulus methods. In the FR scenario, the pair of images are constrained to share the same underlying visual content but differ in types and intensities of distortions. Additionally, a reference image is supplied to facilitate quality comparison. In the NR scenario, the pair of images can be of different content, without supplying any reference image. Upon constructing the pairwise preference matrix $C\in\mathbb{R}^{M\times M}$, where
$C_{ij}=\begin{cases}\#\text{ times }x_{i}\text{ preferred over }x_{j}&\text{if}i\neq j\\ 0&\text{if }i=j,\end{cases}$ (1)
we adopt the maximum a posteriori estimation to aggregate pairwise rankings, under the Thurstone’s Case V assumption [47]. This amounts to solving the following convex optimization problem:
[…]
subject to $\sum_{i}\mu_{i}=0,$
where $\mu=[\mu_{1},\mu_{2},\dots,\mu_{M}]$ are the global quality scores for the $M$ images and $\mathrm{\Phi}(\cdot)$ is the standard normal cumulative distribution function. Alternative pairwise ranking aggregation methods can be employed, such as the Thurstone-Mosteller least squares, the Perron-Frobenius ranking, and the ELO system [15]. We find empirically that these approaches all yield reliable and consistent ranking outcomes. Arguably the double-stimulus method is considered more reliable than the single-stimulus counterpart when gathering human opinions of image quality, despite its $O(M^{2})$ sample complexity. It remains to be seen whether such reliability extends to prompting MLLMs.
Multiple-stimulus Method. As shown in Figure 2 (c), the multiple-stimulus method proposes a more efficient approach to gather partial rankings by simultaneously presenting a set of $L$ images to MLLMs for ranking. Just like the double-stimulus method, in the FR scenario, we limit the list of $L$ images to have identical visual content and supply the corresponding reference image. This restriction and the presence of the reference image are not assumed in the NR scenario. The listwise ranking of $L$ images yields $\binom{L}{2}$ pairwise rankings, which are used to form the pairwise preference matrix $C$ in Eq. (1), followed by ranking aggregation using Eq. (2).
Figure 3 gives the detailed text prompts to implement the single-stimulus, double-stimulus and multiple-stimulus methods for GPT-4V in the NR-IQA scenario. It is crucial to acknowledge that our priority is not to optimize individual textual prompts at the token level, but to explore the impact systematic prompting methods on IQA.
Figure 3: Instantiations of systematic prompting strategies for GPT-4V in the NR scenario. (a) Standard prompting. (b) Chain-of-thought prompting. (c) In-context prompting. See complete FR and NR text prompts in the supplementary material.
3.2 Systematic Prompting from NLP
MLLMs evolve from LLMs, thereby inherently accommodating systematic prompting strategies in NLP.
Standard Prompting. As shown in Figure 3 (a), the standard (and the most basic) prompting method is to directly query the quality score, comparison and ranking results without any exemplars [44] nor descriptions to elicit intermediate reasoning steps. This type of prompting is characterized by its brevity, which has been used in [73] to implement the double-stimulus method.
Chain-of-thought Prompting. As a simple yet powerful approach to encouraging step-by-step reasoning, chain-of-thought prompting [54] has proven to be highly beneficial for a range of NLP tasks. In the same vein, we request MLLMs to detail the perceptual quality of the test image(s) by examining various visual attributes (and comparing to the reference image(s) when available), before integrating the analysis results into overall quality estimate(s) (see Figure 3 (b)).
In-context Prompting. One of the key breakthroughs of LLMs is their ability to perform in-context (i.e., few-shot) prompting [65]. By incorporating just a handful of exemplars during inference, LLMs are adept at making accurate predictions, even on tasks they have not encountered before. Such ability to adapt without additional training is particularly valuable and worth investigating for MLLMs. Thus, as depicted in Figure 3 (c), we apply the in-context prompting method by demonstrating MLLMs image(s) with human quality score(s). Subsequently, we ask MLLMs to assess the quality of the test image(s).
By integrating prompting strategies derived from psychophysics and NLP, we end up with nine prospective prompting systems for testing. Given the variations in model architectures, training data, optimization pipelines, and alignment paradigms [38, 74, 39], it is anticipated that different MLLMs may function optimally with different prompting systems.
3.3 Computational Procedure for Difficult Sample Selection
Inference with MLLMs tends to be slow and costly, making it impractical to evaluate MLLMs on the full IQA datasets paired with the proposed nine prompting systems. We describe a computational procedure to pinpoint a smaller set of the most informative samples that meet three key criteria.
First, they should be difficult, with a high likelihood of causing MLLMs to err. The Group MAximum Differentiation (gMAD) competition [36] has been used before for difficult sample selection, but it is not applicable here due to the violation of affordable inference assumption. We thus utilize efficient expert IQA models as proxies, and identify their failures by maximizing the MSE between model predictions and human quality scores, as a form of “black-box” attacks [40] on MLLMs. Second, the selected samples should be diverse to highlight various aspects of MLLMs’ potential weaknesses in IQA. Third, they should be consistent with small variations in individual quality scores. That is, samples that have inconsistent or uncertain human ratings are less meaningful to assess MLLMs. Putting together, we have the following optimization problem in the FR scenario:
$y_{n}^{\star}=\underset{y\in\mathcal{Y}\setminus\mathcal{Y}^{\star}_{n-1}}{\mathrm{argmax}}\,\frac{1}{|\mathcal{X}_{y}|}\sum_{x\in\mathcal{X}_{y}}\frac{\left(d_{w}(x,y)-q(x)\right)^{2}}{(\sigma(x))^{2}+\epsilon}+\lambda\mathrm{Div}(y,\mathcal{Y}^{\star}_{n-1}),$ (3)
where $\mathcal{Y}^{\star}_{n-1}=\{y^{\star}_{n^{\prime}}\}_{n^{\prime}=1}^{n-1}$ contains the selected reference images in the previous $n-1$ iterations from $\mathcal{Y}$, the set of all reference images. $\mathcal{X}_{y}$ includes all distorted images that originate from $y$. $d_{w}(\cdot,\cdot)$ denotes the expert FR-IQA model, parameterized by $w$. $q(\cdot)$ and $\sigma(\cdot)$ are respectively the human quality score and standard deviation. $\mathrm{Div}(\cdot)$ is the point-to-set measure, quantifying the added diversity of $y$ to $\mathcal{Y}^{\star}_{n-1}$. $\lambda$ is a parameter, trading off the variance-normalized squared error and the diversity measure. $\epsilon$ is a small positive constant to avoid any potential division by zero.
For each of the $N$ identified reference images, we sample top-$K$ difficult distorted images by solving
$x_{k}^{\star}=\underset{x\in\mathcal{X}_{y}\setminus\mathcal{X}^{\star}_{k-1}}{\mathrm{argmax}}\,\frac{(d_{w}(x,y)-q(x))^{2}}{(\sigma(x))^{2}+\epsilon},\quad y\in\mathcal{Y}^{\star}_{N},$ (4)
where $\mathcal{X}^{\star}_{k-1}=\{x^{\star}_{k^{\prime}}\}_{k^{\prime}=1}^{k-1}$. Similarly, in the NR scenario, we sample top-$N$ difficult images by slightly modifying Eq. (3):
$x_{n}^{\star}=\underset{x\in\mathcal{X}\setminus\mathcal{X}^{\star}_{n-1}}{\mathrm{argmax}}\,\frac{\left(q_{w}(x)-q(x)\right)^{2}}{(\sigma(x))^{2}+\epsilon}+\lambda\mathrm{Div}(x,\mathcal{X}^{\star}_{n-1}),$ (5)
where $q_{w}(\cdot)$ denotes the expert NR-IQA model. Figure 4 shows the comparison between difficult sample selection with and without variance normalization (as a measure of sample uncertainty) under the same level of sample diversity. We find that variance-normalized sampling induces a noticeable shift towards zero in the empirical distribution of sample standard deviations, suggesting an enhanced level of human consistency in assessing image quality.
Figure 4: Comparison between difficult sample selection with and without variance normalization under the same level of sample diversity.
4 Experiments
In this section, we commence by presenting the experimental setups, and subsequently undertake a comprehensive evaluation of the nine prompting systems.
4.1 Experimental Setups
Datasets. We examine four visual attributes in the FR scenario, including synthetic structural and textural distortions, texture similarity, color differences, and geometric transformations. The selected datasets are FR-KADID [30], TQD [11], SPCD [51], and Aug-KADID [^2] We follow [11] to introduce mild geometric transformations. [30], respectively. In the NR scenario, we examine synthetic and authentic structural and textual distortions and algorithm-based artifacts, using the NR-KADID [30], SPAQ [16], and AGIQA-3K [25] datasets, respectively.
Sample Selection Details. In the FR scenario, the numbers of sampled reference images $N$ and the corresponding distorted images $K$ are set to $15$ and $10$, respectively. In the NR scenario, the number of selected images $N$ is set to $150$. The trade-off parameter $\lambda$ and the stability constant $\epsilon$ in Eqs. (3) and (5) are set to $0.01$ and $1$, respectively. The point-to-set diversity measure $\mathrm{Div}(\cdot)$ is implemented by the MSE in the feature space of the CLIP visual encoder [42]. The expert FR-IQA model $d_{w}(\cdot,\cdot)$ and the expert NR-IQA model $q_{w}(\cdot)$ are linearly fused with equal weightings using two top-performing models - one excelling in maintaining within-dataset correlation, and the other demonstrating robust cross-dataset generalization. Specifically, for within-dataset performance, we select TOPIQ [8] on FR-KADID [30], DISTS [11] on TQD [11] and Aug-KADID [30], CDNet [51] on SPCD [51], LIQE [71] on SPAQ [16] and NR-KADID [30], and Q-Align [57] on AGIQA-3K [25]. For cross-dataset performance, we choose LPIPS [69] and MANIQA [60] in the FR and NR scenarios, respectively.
Table 1: SRCC results of MLLMs paired with different prompting systems on uniformly sampled images. Model-S, Model-I, and Model-C denote the standard prompting, in-context prompting, and chain-of-thought prompting, respectively. The best result in each section is highlighted in bold.
Method FR IQA NR IQA
FR-KADID Aug-KADID TQD SPCD NR-KADID SPAQ AGIQA-3K
Single-stimulus Method
LLaVA-v1.6-S 0.227 0.013 0.180 0.001 0.262 0.544 0.614
mPLUG-Owl2-S 0.285 0.218 0.228 0.081 0.126 0.467 0.279
InternLM-XC2-VL-S 0.274 0.272 0.299 0.009 0.252 0.794 0.512
GPT-4V-S 0.745 0.786 0.773 0.098 0.467 0.860 0.420
LLaVA-v1.6–I 0.249 0.194 0.222 0.147 0.116 0.019 0.061
mPLUG-Owl2-I 0.373 0.373 0.246 0.047 0.017 0.083 0.409
InternLM-XC2-VL-I 0.380 0.241 0.204 0.087 0.188 0.342 0.461
GPT-4V-I 0.771 0.753 0.738 0.028 0.590 0.845 0.650
LLaVA-v1.6-C 0.164 0.300 0.226 0.174 0.151 0.550 0.580
mPLUG-Owl2-C 0.387 0.361 0.278 0.122 0.179 0.455 0.409
InternLM-XC2-VL-C 0.237 0.306 0.167 0.063 0.306 0.649 0.507
GPT-4V-C 0.809 0.782 0.809 0.121 0.517 0.869 0.677
Double-stimulus Method
LLaVA-v1.6-S 0.387 0.396 0.390 0.113 0.270 0.430 0.234
mPLUG-Owl2-S 0.435 0.307 0.350 0.117 0.126 0.157 0.020
InternLM-XC2-VL-S 0.309 0.408 0.440 0.042 0.267 0.690 0.555
GPT-4V-S 0.679 0.743 0.655 0.031 0.552 0.834 0.599
LLaVA-v1.6-I 0.379 0.396 0.324 0.032 0.169 0.128 0.156
mPLUG-Owl2-I 0.257 0.257 0.169 0.083 0.078 0.164 0.120
InternLM-XC2-VL-I 0.348 0.376 0.379 0.144 0.034 0.108 0.123
GPT-4V-I 0.470 0.244 0.340 0.122 0.531 0.761 0.714
LLaVA-v1.6-C 0.332 0.355 0.257 0.109 0.124 0.065 0.174
mPLUG-Owl2-C 0.409 0.334 0.318 0.013 0.199 0.122 0.130
InternLM-XC2-VL-C 0.332 0.411 0.267 0.131 0.165 0.556 0.546
GPT-4V-C 0.818 0.830 0.786 0.124 0.639 0.881 0.771
Multiple-stimulus Method
LLaVA-v1.6-S 0.349 0.351 0.315 0.241 0.169 0.221 0.210
mPLUG-Owl2-S 0.385 0.428 0.297 0.104 0.124 0.061 0.228
InternLM-XC2-VL-S 0.484 0.420 0.241 0.015 0.047 0.044 0.154
GPT-4V-S 0.824 0.844 0.747 0.037 0.397 0.715 0.461
LLaVA-v1.6-I 0.337 0.380 0.356 0.203 0.152 0.033 0.241
mPLUG-Owl2-I 0.268 0.268 0.377 0.067 0.196 0.142 0.121
InternLM-XC2-VL-I 0.489 0.235 0.212 0.046 0.038 0.102 0.114
GPT-4V-I 0.585 0.496 0.389 0.023 0.168 0.416 0.201
LLaVA-v1.6-C 0.292 0.424 0.288 0.043 0.227 0.111 0.122
mPLUG-Owl2-C 0.377 0.406 0.376 0.126 0.214 0.166 0.084
InternLM-XC2-VL-C 0.500 0.466 0.273 0.038 0.031 0.037 0.148
GPT-4V-C 0.761 0.806 0.754 0.036 0.537 0.817 0.679
Figure 5: Behaviors of different MLLMs in recognizing objects from multiple images.
4.2 Comparison of Nine Prompting Systems
We initially combine the nine prompting systems with three open-source MLLMs: LLaVA-v1.6 (Mistral-7B) [31], InternLM-XComposer2-VL (InernLM2-7B) [13] and mPLUG-Owl2 (LLaMA2-7B) [64] and one close-source MLLM: GPT-4V [61], and compare them on $150$ uniformly sampled images, with the goal of identifying the best prompting system for each MLLM. Table 1 shows the Spearman’s rank correlation coefficient (SRCC) results.
Analysis of the Two IQA Scenarios. From the Table, it is evident that no open-source MLLMs achieve satisfactory IQA results in the FR scenario regardless of the prompting system used. These models may generate irrelevant text outputs or encounter complete failures when the input text prompts become more detailed and elaborate. We hypothesize that these models are predominantly trained or aligned on single-image vision tasks, making them struggle with analyzing multiple images, especially of the same underlying content [31]. Figure 5 provides a simple example of how different MLLMs behave when recognizing objects from multiple images, a high-level vision task at which they should excel. However, it appears that LLaVA-v1.6 [31] and mPLUG-Owl2 [64] completely disregard the second image, despite it being explicitly mentioned and separated from the first image in the input prompt. In the NR scenario with single-stimulus standard prompting (i.e., the single-image analysis scenario), LLaVA-v1.6 [31] and mPLUG-Owl2 [64] deliver improved quality prediction accuracy. Moreover, InternLM-XComposer2-VL [13] and GPT-4V perform remarkably in handling realistic camera distortions on SPAQ [16].
Analysis of Psychophysical Prompting Methods. The results in the table reveal that for the three open-source MLLMs, the single-stimulus method is the optimal choice due to their limited ability to analyze multiple images. In stark contrast, GPT-4V [61] benefits from multiple-image analysis, and performs optimally under the double-stimulus chain-of-thought prompting. These findings raise questions about recent claims [58] that open-source MLLMs have surpassed GPT-4V-level IQA performance by instruction tuning on existing (weakly-)annotated benchmarks, given that the optimal prompting system has not been paired with GPT-4V.
Figure 6: Comparison of the double-stimulus standard and chain-of-thought promptings for GPT-4V in the NR scenario.
Analysis of NLP Prompting Methods. It is expected that in-context prompting is little likely to bring performance gains to the three open-source MLLMs due to the added complexity of processing additional image(s) in the context. Interestingly, we find that in-context prompting does not aid GPT-4V with the single-stimulus prompting. When combined with the multiple-stimulus prompting, GPT-4V performs poorly across various visual attributes and in both the FR and NR scenarios. To take a closer look, we compare the single-stimulus in-context prompting with double-stimulus standard prompting for GPT-4V. The main difference is that the former offers a human quality score for the contextual image, aiming to aid visual quality comparison. The comparable performance achieved by both prompting systems suggests that the inclusion of a human quality score has minimal impact on the GPT-4V inference. In cases when the input prompt contains a large number of images (e.g., eight images for the multiple-stimulus in-context prompting), even GPT-4V finds difficult in processing such substantial amount of visual information, resulting in a sharp drop in prediction accuracy.
The chain-of-thought prompting, on the other hand, paints a different picture: it consistently enhances the performance of GPT-4V [61] in conjunction with three psychophysical testing protocols and across nearly all visual attributes. The remarkable improvements may arise because the chain-of-thought prompting elicits the break down of the intricate IQA task into simpler sub-tasks. This allows for a more meticulous inspection of visually critical factors such as structure and texture preservation, as well as color and luminance reproduction. On the other hand, as shown in Figure 6, the chain-of-thought prompting encourages absorbing IQA into a boarder and higher-level task, orchestrating with additional physical, geometrical, and semantic information of the natural scene to infer image quality. This reminds the authors of a famous slogan in computer vision: “If you cannot solve a simple problem in vision, you may have to solve a complex one (by Songchun Zhu)”.
Table 2: Comparison of MLLMs with optimally suited prompting systems against expert IQA systems in the FR scenario. Black and blue numbers in bold represent the best expert IQA systems and the MLLMs, respectively. * indicates that the model has been trained on the dataset.
FR-KADID Aug-KADID TQD SPCD
Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
PSNR 0.479 0.675 0.381 0.644 0.345 0.522 0.576 0.570
SSIM [52] 0.553 0.694 0.405 0.633 0.510 0.618 0.229 0.246
FSIM [68] 0.704 0.762 0.400 0.560 0.332 0.408 0.205 0.206
LPIPS [69] 0.477 0.654 0.547 0.654 0.469 0.511 0.280 0.252
AHIQ [24] 0.512 0.583 0.512 0.688 0.467 0.608 0.240 0.269
DISTS [11] 0.647* 0.740* 0.701 0.696 0.911 0.901 0.454 0.422
LLaVA-v1.6 [31] 0.112 0.218 0.198 0.213 0.180 0.226 0.037 0.008
mPLUG-Owl2 [64] 0.248 0.435 0.358 0.484 0.228 0.335 0.102 0.108
InternLM-XC2-VL [13] 0.246 0.336 0.235 0.404 0.299 0.421 0.171 0.143
GPT-4V [61] 0.669 0.795 0.708 0.800 0.786 0.857 0.122 0.234
4.3 Further Testing on Difficult Data
In this subsection, we compare the previous four MLLMs, each with the optimally suited prompting system and a quality-instruction-tuned MLLM, Q-Instruct [56], against representative expert IQA systems, including PSNR, SSIM [52], FSIM [68], LPIPS [69], AHIQ [24] and DISTS [11] as FR models, and NIQE [37], MUSIQ [21], MANIQA [60], LIQE [71] as NR models on the set of difficult data described in Sec. 3.3.
Results in the FR Scenario. Table 2 shows the SRCC and Pearson’s linear correlation coefficient (PLCC) results in the FR scenario. The primary observation is that the selected difficult images by “attacking” expert IQA models using Eqs. (3) and (5) pose a challenge to GPT-4V, with reduced IQA performance. Nevertheless, on FR-KADID with synthetic structural and textural distortions, GPT-4V [61] is on par with the leading expert IQA model, FSIM [68], which assesses structural similarity in the feature space. On Aug-KADID and TQD with geometric transformations and texture similarity, respectively, GPT-4V exhibits a commendable level of resilience, although not as good as DISTS [11]. It is important to recognize that the ways of attaining these aspects of perceptual robustness differ between the two methods. DISTS relies on comparison of spatial averages of VGG [45] feature maps as a complete and holistic texture characterization, whereas GPT-4V benefits from its intrinsic ability to compile and process data sequentially (as highlighted in the Introduction). On SPCD with color differences, all MLLMs including GPT-4V encounter challenges in emulating human color perception, and struggle to differentiate images even with clearly noticeable variations in color appearances (to the human eye).
Table 3: Comparison of MLLMs with optimally suited prompting systems against expert IQA systems in the NR scenario.
SPAQ NR-KADID AGIQA-3K
Method SRCC PLCC SRCC PLCC SRCC PLCC
NIQE [37] 0.551 0.616 0.385 0.555 0.610 0.651
MUSIQ [21] 0.769 0.817 0.567 0.653 0.686 0.588
MANIQA [60] 0.546 0.564 0.428 0.387 0.521 0.599
LIQE [71] 0.781* 0.752* 0.866* 0.930* 0.703 0.693
LLaVA-v1.6 [31] 0.317 0.305 0.428 0.370 0.503 0.573
mPLUG-Owl2 [64] 0.270 0.198 0.128 0.187 0.168 0.201
InternLM-XC2-VL [13] 0.580 0.540 0.454 0.361 0.608 0.590
Q-Instruct [56] 0.799* 0.783* 0.635 0.613 0.853* 0.821*
GPT-4V [61] 0.845 0.843 0.513 0.453 0.783 0.746
Results in the NR Scenario. Table 3 shows the SRCC and PLCC results in the NR scenario. It is clear that GPT-4V [61] is outstanding in capturing the authentic structural and textural distortions on SPAQ, surpassing the two expert IQA systems MUSIQ [21] and LIQE [71]. Notably, the latter has prior exposure to the SPAQ dataset. Furthermore, GPT-4V showcases a remarkable generalizablity to AI-generated images, indicating its great potential to guide the optimization of generative AI models for images. Q-Instruct [56], a model fine-tuned from a variant of LLaVA, exhibits significantly enhanced IQA skills compared to LLaVA-v1.6, thus highlighting the effectiveness of visual instruction tuning (particularly on images with similar distortion appearances).
4.4 Discussion and Limitation
We briefly discuss some of the limitations of this work and opportunities for future work. First, the current prompting systems have room for improvement, as the input prompts are not optimized. This presents an opportunity for exploring automatic prompt optimization [27, 18] within our prompting systems. Second, our sampler relies on human quality scores, and is thus only applicable to existing labeled IQA datasets. Extending it to sample from large-scale sets of unlabeled images will need to 1) accelerate the inference speed of MLLMs so as to leverage model falsification methodologies (e.g., the gMAD competition [36]) or 2) train additional failure-prediction modules [7] in a parameter-efficient way [19]. Third, the textual responses produced by MLLMs have not been quantitatively assessed. Prior research has investigated the application of GPT-4 [31] to assess these responses in terms of correctness, consistency, relevance, informativeness, coherence, and naturalness. Given our focus on responses pertinent to visual quality, it seems reasonable to develop a semantic textual similarity model, specifically designed for the IQA task, with improved accuracy and robustness. Fourth, this work does not touch on instruction tuning of MLLMs to enhance the IQA performance. Preliminary efforts have been made by Wu et al. [56] and You et al. [67] to directly fine-tune open-source MLLMs on datasets with image quality descriptions. Our results suggest a more encouraging approach: incorporating IQA in a boarder vision task for joint optimization, with the goal of achieving an ideal balance between the generality and specificity of pre-trained MLLMs [61, 33, 23].
5 Conclusion
We have presented a comprehensive study of MLLMs for IQA, with emphasis on systematic prompting strategies. Our study arises as a natural combination of methods from two separate lines of research: psychophysics and NLP. The first involves collecting reliable measurements of perceptual quantities in response to physical stimuli. The second endeavours to design input textual descriptions to an LLM to elicit a specific response. Our experiments have shown that different MLLMs admit different prompting systems to work optimally. This suggests the need for a systematic re-evaluation of the recent progress in MLLMs for IQA, particular in comparison to the state-of-the-art GPT-4V [61]. Moreover, there is still ample room in improving MLLMs (including GPT-4V) for IQA, especially fine-grained quality discrimination and multiple-image quality analysis.
Meanwhile, we have emphasized the importance of sample selection when evaluating MLLMs for IQA, owing to the high cost associated with inference. In response, we have proposed a computational procedure for difficult sample selection, taking into account both sample diversity and uncertainty. Our sampler relies on an inference-efficient expert IQA model, and can be seen as a black box attack on MLLMs, assuming no knowledge of their internal mechanisms nor the external input-output behaviors.

### v2

@@
-
+A Comprehensive Study of Multimodal Large Language Models for Image Quality Assessment
 Abstract
-While Multimodal Large Language Models (MLLMs) have experienced significant advancement on visual understanding and reasoning, their potentials to serve as powerful, flexible, interpretable, and text-driven models for Image Quality Assessment (IQA) remains largely unexplored.
+While Multimodal Large Language Models (MLLMs) have experienced significant advancement in visual understanding and reasoning, their potential to serve as powerful, flexible, interpretable, and text-driven models for Image Quality Assessment (IQA) remains largely unexplored.
 In this paper, we conduct a comprehensive and systematic study of prompting MLLMs for IQA.
-Specifically, we first investigate nine prompting systems for MLLMs as the combinations of three standardized testing procedures in psychophysics (i.e., the single-stimulus, double-stimulus, and multiple-stimulus methods) and three popular prompting strategies in natural language processing (i.e., the standard, in-context, and chain-of-thought prompting).
+We first investigate nine prompting systems for MLLMs as the combinations of three standardized testing procedures in psychophysics (i.e., the single-stimulus, double-stimulus, and multiple-stimulus methods) and three popular prompting strategies in natural language processing (i.e., the standard, in-context, and chain-of-thought prompting).
 We then present a difficult sample selection procedure, taking into account sample diversity and uncertainty, to further challenge MLLMs equipped with the respective optimal prompting systems.
-We assess three open-source and one close-source MLLMs on several visual attributes of image quality (e.g., structural and textural distortions, color differences, and geometric transformations) in both full-reference and no-reference scenarios.
-Experimental results show that only the close-source GPT-4V provides a reasonable account for human perception of image quality, but is weak at discriminating fine-grained quality variations (e.g., color differences) and at comparing visual quality of multiple images, tasks humans can perform effortlessly.
+We assess three open-source and one closed-source MLLMs on several visual attributes of image quality (e.g., structural and textural distortions, geometric transformations, and color differences) in both full-reference and no-reference scenarios.
+Experimental results show that only the closed-source GPT-4V provides a reasonable account for human perception of image quality, but is weak at discriminating fine-grained quality variations (e.g., color differences) and at comparing visual quality of multiple images, tasks humans can perform effortlessly.
 Keywords:
-Image quality assessment Multimodal large language models Model benchmarking
-Figure 1: Illustration of several visual attributes of image quality in our experiments.
+Image quality assessment Multimodal large language models Model comparison
+Figure 1: Illustration of visual attributes of image quality in our experiments.
 1 Introduction
 The evolution of Large Language Models (LLMs) has marked a significant milestone in the field of Artificial Intelligence (AI) [51, 47, 6, 2].
-The underlying idea of scaling the model size and training data [20] has rendered LLMs the abilities to perform various Natural Language Processing (NLP) tasks with unprecedented levels of accuracy [2].
+The underlying idea of scaling the model size and training data [19] has rendered LLMs the abilities to perform various Natural Language Processing (NLP) tasks with unprecedented levels of accuracy.
 In the midst of these developments, a particularly promising offshoot has emerged in the form of Multimodal LLMs (MLLMs).
 These advanced models have taken the capabilities of LLMs a step further by incorporating visual data alongside text [29, 62, 13, 4, 43, 59].
-MLLMs typically integrate visual data via Vision Transformers (ViTs) [14] for feature extraction, attention mechanisms [50] for visual-textual relationship modeling, and fusion modules [3, 26, 17] to merge the two modalities.
+MLLMs typically integrate visual data via Vision Transformers (ViTs) [14] for feature extraction, attention mechanisms [48] for visual-textual relationship modeling, and connector modules [3, 24, 16, 45] to merge the two modalities.
 These techniques enable MLLMs to process both text and image data holistically, extending the application scenario of LLMs.
@@
   • Improved robustness.
-Being sequential in nature, MLLMs do not rely on alignment of images for quality assessment, making them easily robust to perturbations that are imperceptible to the human eye, such as mild geometrical transformations [34] and texture substitution [11].
+Being sequential in nature, MLLMs do not rely on the alignment of images for quality assessment, making them easily robust to perturbations that are imperceptible to the human eye, such as mild geometrical transformations [31] and texture resampling [11].
   • Flexibility.
@@
 MLLMs generate descriptive text rather than merely providing a numerical score [66].
-This text output allows for more detailed, contextually rich, and human-like quality assessment, making MLLMs valuable for IQA model-in-the-loop image enhancement [29].
+This text output allows for more detailed, contextually rich, and human-like quality assessment, making MLLMs valuable for IQA model-in-the-loop image processing [27].
 Previous work [53, 54, 66] focuses primarily on establishing IQA datasets with human quality descriptions to benchmark MLLMs in terms of quality question answering, rating, and reasoning.
@@
 After all, the field of IQA has a rich history, and numerous established human-rated image quality datasets are readily accessible for evaluating this perceptual aspect of MLLMs.
-Closest to ours, Zhu et al.
-[73] tested the pairwise quality comparison capability of MLLMs in a constrained context.
-In this work, we further pursue this direction, and conduct a comprehensive and systematic study of prompting MLLMs for IQA, as opposed to optimizing individual text prompts.
+In this work, we conduct a comprehensive and systematic study of prompting MLLMs for IQA.
 We first explore nine prompting systems for MLLMs, combining standardized testing procedures in psychophysics (i.e., the single-stimulus, double-stimulus, and multiple-stimulus methods) with popular prompting strategies in NLP (i.e., the standard, in-context, and chain-of-thought prompting).
-Interestingly, we discover that the optimal prompting system varies between open-source [31, 64, 13] and closed-source MLLMs.
 To further challenge MLLMs, we propose a computational procedure to select difficult samples using top-performing IQA expert models as proxies [58, 70, 11, 68, 55, 8], while taking into account sample diversity and uncertainty.
-Under both Full-Reference (FR) and No-Reference (NR) settings, we experiment with three open-source and one close-source MLLMs across several visual attributes of image quality, including structural and textural distortions, color differences, and geometric transformations (see Fig.
+Under both Full-Reference (FR) and No-Reference (NR) settings, we experiment with three open-source and one closed-source MLLMs across several visual attributes of image quality, including structural and textural distortions, geometric transformations, and color differences (see Fig.
 1).
-Experimental results show that even the most advanced MLLM (i.e., the proprietary GPT-4V) still has room for improvement in modeling the incredibly sophisticated adaptable HVS for IQA.
+Our experimental results highlight three key takeaways.
+First, different MLLMs require different prompting systems to perform optimally.
+This implies the need for a re-examination of the recent progress in MLLMs for IQA, particularly in comparison to GPT-4V [56, 66, 65].
+Second, aided by the proposed difficult sample selection method, we demonstrate that there is still ample room for improving MLLMs (including GPT-4V) for IQA, especially in fine-grained quality discrimination and multiple-image quality analysis.
+Third, we argue that directly fine-tuning open-source MLLMs on datasets with image quality descriptions may not be effective due to the risk of catastrophic forgetting of models’ general abilities in visual reasoning and understanding.
+The encouraging results from the chain-of-thought prompting suggest that IQA should be integrated into a broader and higher-level task, making use of additional physical, geometrical, and semantic information to infer image quality.
 2 Related Work
@@
 2.1 Expert Models for IQA
-IQA can be roughly divided into two categories: FR- and NR-IQA.
-FR-IQA models are preferred in situations where the unimpaired reference image is available, such as in many low-level vision tasks.
-Representative design philosophies for FR-IQA range from measuring error visibility (e.g., the mean squared error (MSE)), structural similarity (e.g., the SSIM index [52]), and mutual information (e.g., the VIF measure [43]) to (deep) learning-based methods (e.g., the LPIPS [69] and DISTS metrics [11]) and fusion-based approaches (e.g., VMAF [48]).
+IQA can be divided into two categories: FR- and NR-IQA.
+FR-IQA models are preferred in situations where the undistorted reference image is available.
+Representative design philosophies for FR-IQA range from measuring error visibility (e.g., the mean squared error (MSE)), structural similarity (e.g., the SSIM index [50]), and mutual information (e.g., the VIF measure [40]) to (deep) learning-based methods (e.g., the LPIPS [68] and DISTS metrics [11]) and to fusion-based approaches (e.g., VMAF [46]).
 Almost all FR-IQA models depend on the proper alignment of the reference and test images to execute co-located comparison at the pixel, patch, or feature level.
 Consequently, these models may struggle to capture the robustness of the HVS to mild geometric transformations and texture resampling.
-Additionally, these models frequently disregard or give a superficial treatment of color information, leading to a poor account for perceptual color differences [51, 9].
-NR-IQA models [37, 70, 60, 21, 71, 59, 63] are the most challenging yet widely applicable as they are designed to evaluate image quality without any reference.
-Regardless of whether they are knowledge-driven [37] or data-driven [70, 71], NR-IQA models rely heavily on the training data, which often demonstrate weak generalization when faced with a broader type of images and distortions.
+Additionally, these models often give a superficial treatment of color information, leading to a poor account for perceptual color differences [49, 9].
+NR-IQA models [34, 69, 58, 20, 70, 57, 61] are more challenging yet widely applicable as they evaluate image quality without any reference.
+Whether they are knowledge-driven [34] or data-driven [69, 70], NR-IQA models rely heavily on human-annotated training data.
+This reliance leads to weak generalization when applied to a wider range of images and distortions.
 Apart from accuracy and generalization, expert IQA models also fall short in terms of flexibility and interpretability.
-This is primarily because they reduce the evaluation of image quality to a numerical score, failing to leverage other formats of inputs, and output other detailed and perceptually relevant information.
+This is primarily because they summarize image quality using a numerical score, missing the opportunity to leverage other input formats and to provide more detailed and perceptually relevant outputs.
 2.2 MLLMs for IQA
-Previous work has been centered on benchmarking and refining MLLMs for IQA.
+Previous work has been centered on benchmarking and fine-tuning MLLMs for IQA.
 The pioneering datasets, DepictQA [66] and Q-Bench [53], have initiated the assessment of MLLMs in quality rating and reasoning under the FR and NR conditions, respectively.
-DepictQA employs a double-stimulus method, prompting MLLMs to compare a pair of images against a standard reference.
-In contrast, Q-Bench takes a single-stimulus approach, assigning a numerical score based on the classification of image quality as “poor” or “good”.
-Notably, these investigations reveal that, except for the proprietary GPT-4V [61], open-source MLLMs like LLaVA [31], Kosmos-2 [41], and MiniGPT-4 [72] show limited success in replicating human perception of image quality.
-Efforts to enhance open-source MLLMs through instruction tuning have been made [56], focusing on the prediction of subsequent tokens.
+DepictQA employs a double-stimulus method, prompting MLLMs to compare a pair of images against the reference.
+Q-Bench takes a single-stimulus approach, computing a numerical score based on the classification of image quality as “poor” or “good”.
+These investigations reveal that, except for the proprietary GPT-4V [59], open-source MLLMs like LLaVA [29], Kosmos-2 [38], and MiniGPT-4 [71] show limited success in replicating human perception of image quality.
+Efforts to enhance open-source MLLMs through instruction tuning have been made [54].
 Yet, such models often revert to generating template-like quality descriptions, lacking the level of flexibility we are looking for.
-In this work, we take a step back and conduct a comprehensive and systematic exploration of prompting techniques for MLLMs on existing human-rated IQA datasets.
-Figure 2: Three standardized psychophysical testing procedures for IQA. (a) Single-stimulus method. (b) Double-stimulus method. (c) Multiple-stimulus method.
-3 Systematic Prompting MLLMs for IQA
-In this section, we introduce several prompting systems for MLLMs in the context of IQA, followed by describing the difficult sample selection procedure.
-3.1 Systematic Prompting from Psychophysics
+In this work, we take a step back and conduct a comprehensive and systematic exploration of prompting techniques for MLLMs on existing human-rated image quality datasets.
+3 Prompting MLLMs for IQA
+In this section, we introduce several IQA prompting systems for MLLMs, followed by the description of the difficult sample selection procedure.
+3.1 Prompting Strategies from Psychophysics
 Although MLLMs and the HVS (or more generally the human brain) operate in different domains, they have fascinating similarities, especially in representational hierarchy [5], visual-textual integration, contextual reasoning [21], and adaptation [12].
-Thus, it is natural to employ standardized psychophysical testing procedures as the systematic prompting methods for IQA.
+Thus, it is natural to employ standardized psychophysical testing procedures as prompting strategies for IQA.
 Single-stimulus Method.
 As depicted in Figure 2 (a), given a test image $x\in\mathcal{X}$, the FR single-stimulus method [^1] This is also known as the double-stimulus impairment rating by treating the reference image as a second visual stimulus.
-suggests MLLMs to output a quality score, $q(x)\in[0,100]$, based on its perceptual difference to the corresponding reference image $y\in\mathcal{Y}$.
-Here, a larger number indicates higher quality.
-In contrast, the NR single-stimulus method derives the quality score solely from the perceptually important visual attributes of the test image $x$.
-Generally, the single-stimulus method is scalable and straightforward to implement for MLLMs, but may not capture relative quality accurately and may exhibit context dependency (e.g., the presentation order).
+suggests MLLMs to output a quality score, $q(x)\in[0,100]$, based on its perceptual distance to the corresponding reference image $y\in\mathcal{Y}$.
+Here, a larger $q$ indicates higher quality.
+The NR single-stimulus method derives the quality score solely from the test image $x$.
+Generally, the single-stimulus method is scalable and straightforward to implement for MLLMs, but may not capture relative quality accurately.
+Figure 2: Three standardized psychophysical testing procedures for IQA. (a) Single-stimulus method. (b) Double-stimulus method. (c) Multiple-stimulus method.
 Double-stimulus Method.
-Also known as two-alternative forced choice (2AFC) and paired comparison, the double-stimulus method first samples a pair of images either uniformly or actively [62] from a set of images $\mathcal{X}=\{x_{i}\}^{M}_{i=1}$, where $M$ is the number of images.
+Also known as two-alternative forced choice (2AFC) and paired comparison, the double-stimulus method first samples a pair of images either uniformly or actively [60] from a set of images $\mathcal{X}=\{x_{i}\}^{M}_{i=1}$, where $M$ is the total number of images.
 In our implementation, MLLMs are not forced to select between the two alternatives with higher quality.
-Instead, they have a third option: the quality of the two images is comparable.
+Instead, they have a third option to indicate that the quality of the two images is comparable.
 Figure 2 (b) illustrates the FR and NR double-stimulus methods.
-In the FR scenario, the pair of images are constrained to share the same underlying visual content but differ in types and intensities of distortions.
-Additionally, a reference image is supplied to facilitate quality comparison.
+In the FR scenario, the pair of images are constrained to share the same underlying visual content but differ in distortion types and intensities.
+The reference image is accompanied to facilitate quality comparison.
 In the NR scenario, the pair of images can be of different content, without supplying any reference image.
-Upon constructing the pairwise preference matrix $C\in\mathbb{R}^{M\times M}$, where
-$C_{ij}=\begin{cases}\#\text{ times }x_{i}\text{ preferred over }x_{j}&\text{if}i\neq j\\ 0&\text{if }i=j,\end{cases}$ (1)
-we adopt the maximum a posteriori estimation to aggregate pairwise rankings, under the Thurstone’s Case V assumption [47]. This amounts to solving the following convex optimization problem:
-[…]
-subject to $\sum_{i}\mu_{i}=0,$
-where $\mu=[\mu_{1},\mu_{2},\dots,\mu_{M}]$ are the global quality scores for the $M$ images and $\mathrm{\Phi}(\cdot)$ is the standard normal cumulative distribution function.
-Alternative pairwise ranking aggregation methods can be employed, such as the Thurstone-Mosteller least squares, the Perron-Frobenius ranking, and the ELO system [15].
-We find empirically that these approaches all yield reliable and consistent ranking outcomes.
-Arguably the double-stimulus method is considered more reliable than the single-stimulus counterpart when gathering human opinions of image quality, despite its $O(M^{2})$ sample complexity.
+Upon constructing the pairwise preference matrix $C\in\mathbb{R}^{M\times M}$, where $C_{ij}$ records the number of times $x_{i}$ is preferred over $x_{j}$ for $i\neq j$, we adopt the maximum a posteriori estimation to aggregate pairwise rankings under the Thurstone’s Case V model [44].
+Arguably, the double-stimulus method is considered more reliable than the single-stimulus one when gathering human opinions of image quality, despite its $O(M^{2})$ sample complexity.
 It remains to be seen whether such reliability extends to prompting MLLMs.
 Multiple-stimulus Method.
-As shown in Figure 2 (c), the multiple-stimulus method proposes a more efficient approach to gather partial rankings by simultaneously presenting a set of $L$ images to MLLMs for ranking.
+As shown in Figure 2 (c), the multiple-stimulus method presents a more efficient approach to gathering partial rankings by simultaneously presenting a set of $L$ images to MLLMs.
 Just like the double-stimulus method, in the FR scenario, we limit the list of $L$ images to have identical visual content and supply the corresponding reference image.
 This restriction and the presence of the reference image are not assumed in the NR scenario.
-The listwise ranking of $L$ images yields $\binom{L}{2}$ pairwise rankings, which are used to form the pairwise preference matrix $C$ in Eq.
-(1), followed by ranking aggregation using Eq.
-(2).
-Figure 3 gives the detailed text prompts to implement the single-stimulus, double-stimulus and multiple-stimulus methods for GPT-4V in the NR-IQA scenario.
-It is crucial to acknowledge that our priority is not to optimize individual textual prompts at the token level, but to explore the impact systematic prompting methods on IQA.
+The listwise ranking of $L$ images yields $\binom{L}{2}$ pairwise rankings, which are used to form the pairwise preference matrix $C$, followed by ranking aggregation.
+Figure 3 gives the detailed text prompts to implement the single-stimulus, double-stimulus, and multiple-stimulus methods for GPT-4V in the NR-IQA scenario.
+3.2 Prompting Strategies from NLP
+Standard Prompting.
+As shown in Figure 3 (a), the standard (and the most basic) prompting method is to directly query the quality score, comparison and ranking results without any exemplars [41] nor descriptions to elicit intermediate reasoning steps.
+This type of prompting is characterized by its brevity, which has been used in [72] to implement the double-stimulus method.
 Figure 3: Instantiations of systematic prompting strategies for GPT-4V in the NR scenario.
@@
 See complete FR and NR text prompts in the supplementary material.
-3.2 Systematic Prompting from NLP
-MLLMs evolve from LLMs, thereby inherently accommodating systematic prompting strategies in NLP.
-Standard Prompting.
-As shown in Figure 3 (a), the standard (and the most basic) prompting method is to directly query the quality score, comparison and ranking results without any exemplars [44] nor descriptions to elicit intermediate reasoning steps.
-This type of prompting is characterized by its brevity, which has been used in [73] to implement the double-stimulus method.
 Chain-of-thought Prompting.
-As a simple yet powerful approach to encouraging step-by-step reasoning, chain-of-thought prompting [54] has proven to be highly beneficial for a range of NLP tasks.
-In the same vein, we request MLLMs to detail the perceptual quality of the test image(s) by examining various visual attributes (and comparing to the reference image(s) when available), before integrating the analysis results into overall quality estimate(s) (see Figure 3 (b)).
+As a simple yet powerful approach to encouraging step-by-step reasoning, chain-of-thought prompting [52] has proven beneficial for a range of NLP tasks.
+In the same vein, we request MLLMs to detail the perceptual quality of the test image(s) by examining various visual attributes (and comparing them to the reference image(s) when available), before integrating the analysis results into the overall quality estimate(s).
 In-context Prompting.
@@
 Subsequently, we ask MLLMs to assess the quality of the test image(s).
-By integrating prompting strategies derived from psychophysics and NLP, we end up with nine prospective prompting systems for testing.
-Given the variations in model architectures, training data, optimization pipelines, and alignment paradigms [38, 74, 39], it is anticipated that different MLLMs may function optimally with different prompting systems.
+By integrating prompting strategies derived from psychophysics and NLP, we end up with nine candidate prompting systems.
+Given the variations in training data, model architectures, optimization pipelines, and alignment strategies [35, 73, 36], it is anticipated that different MLLMs may function optimally with different prompting systems.
 3.3 Computational Procedure for Difficult Sample Selection
-Inference with MLLMs tends to be slow and costly, making it impractical to evaluate MLLMs on the full IQA datasets paired with the proposed nine prompting systems.
-We describe a computational procedure to pinpoint a smaller set of the most informative samples that meet three key criteria.
+Inference with MLLMs tends to be slow and costly, making it impractical to evaluate MLLMs on the full IQA datasets, each paired with the nine candidate prompting systems.
+We describe a computational procedure to pinpoint a smaller set of the most informative testing samples with three desired properties.
 First, they should be difficult, with a high likelihood of causing MLLMs to err.
-The Group MAximum Differentiation (gMAD) competition [36] has been used before for difficult sample selection, but it is not applicable here due to the violation of affordable inference assumption.
-We thus utilize efficient expert IQA models as proxies, and identify their failures by maximizing the MSE between model predictions and human quality scores, as a form of “black-box” attacks [40] on MLLMs.
+The Group MAximum Differentiation (gMAD) competition [33] has been used before for difficult sample selection, but it is not applicable here due to the violation of the affordable inference assumption.
+We thus utilize efficient expert IQA models as proxies, and identify their failures by maximizing the MSE between model predictions and human quality scores as a form of “black-box” attacks [37] on MLLMs.
 Second, the selected samples should be diverse to highlight various aspects of MLLMs’ potential weaknesses in IQA.
-Third, they should be consistent with small variations in individual quality scores.
-That is, samples that have inconsistent or uncertain human ratings are less meaningful to assess MLLMs.
-Putting together, we have the following optimization problem in the FR scenario:
+Third, they should be consistent with the majority of human judgments with small variations.
+Putting together, we formulate the following optimization problem in the FR scenario:
 $y_{n}^{\star}=\underset{y\in\mathcal{Y}\setminus\mathcal{Y}^{\star}_{n-1}}{\mathrm{argmax}}\,\frac{1}{|\mathcal{X}_{y}|}\sum_{x\in\mathcal{X}_{y}}\frac{\left(d_{w}(x,y)-q(x)\right)^{2}}{(\sigma(x))^{2}+\epsilon}+\lambda\mathrm{Div}(y,\mathcal{Y}^{\star}_{n-1}),$ (1)
@@
 $\mathcal{X}_{y}$ includes all distorted images that originate from $y$.
-$d_{w}(\cdot,\cdot)$ denotes the expert FR-IQA model, parameterized by $w$.
+$d_{w}(\cdot,\cdot)$ denotes a generic expert FR-IQA model, parameterized by $w$.
 $q(\cdot)$ and $\sigma(\cdot)$ are respectively the human quality score and standard deviation.
-$\mathrm{Div}(\cdot)$ is the point-to-set measure, quantifying the added diversity of $y$ to $\mathcal{Y}^{\star}_{n-1}$.
+$\mathrm{Div}(\cdot)$ is a point-to-set distance measure, quantifying the added diversity of $y$ to $\mathcal{Y}^{\star}_{n-1}$.
 $\lambda$ is a parameter, trading off the variance-normalized squared error and the diversity measure.
 $\epsilon$ is a small positive constant to avoid any potential division by zero.
+Figure 4: Comparison between difficult sample selection with and without variance normalization under the same level of sample diversity.
 For each of the $N$ identified reference images, we sample top-$K$ difficult distorted images by solving
@@
 $x_{n}^{\star}=\underset{x\in\mathcal{X}\setminus\mathcal{X}^{\star}_{n-1}}{\mathrm{argmax}}\,\frac{\left(q_{w}(x)-q(x)\right)^{2}}{(\sigma(x))^{2}+\epsilon}+\lambda\mathrm{Div}(x,\mathcal{X}^{\star}_{n-1}),$ (3)
-where $q_{w}(\cdot)$ denotes the expert NR-IQA model.
+where $q_{w}(\cdot)$ denotes a generic expert NR-IQA model.
 Figure 4 shows the comparison between difficult sample selection with and without variance normalization (as a measure of sample uncertainty) under the same level of sample diversity.
 We find that variance-normalized sampling induces a noticeable shift towards zero in the empirical distribution of sample standard deviations, suggesting an enhanced level of human consistency in assessing image quality.
-Figure 4: Comparison between difficult sample selection with and without variance normalization under the same level of sample diversity.
 4 Experiments
-In this section, we commence by presenting the experimental setups, and subsequently undertake a comprehensive evaluation of the nine prompting systems.
+In this section, we first present the experimental setups, and then carry out a comprehensive evaluation of the nine IQA prompting systems for MLLMs.
 4.1 Experimental Setups
 Datasets.
-We examine four visual attributes in the FR scenario, including synthetic structural and textural distortions, texture similarity, color differences, and geometric transformations.
+We examine four visual attributes in the FR scenario, including synthetic structural and textural distortions, geometric transformations, texture similarity, and color differences.
 The selected datasets are FR-KADID [28], TQD [11], SPCD [49], and Aug-KADID [^2] We follow [11] to introduce mild geometric transformations.
@@
 (1) and (3) are set to $0.01$ and $1$, respectively.
-The point-to-set diversity measure $\mathrm{Div}(\cdot)$ is implemented by the MSE in the feature space of the CLIP visual encoder [42].
-The expert FR-IQA model $d_{w}(\cdot,\cdot)$ and the expert NR-IQA model $q_{w}(\cdot)$ are linearly fused with equal weightings using two top-performing models - one excelling in maintaining within-dataset correlation, and the other demonstrating robust cross-dataset generalization.
+The point-to-set distance measure $\mathrm{Div}(\cdot)$ is implemented by the MSE in the feature space of the CLIP visual encoder [39].
+The expert FR-IQA model $d_{w}(\cdot,\cdot)$ and NR-IQA model $q_{w}(\cdot)$ are linearly fused with equal weightings using two top-performing models - one excelling in maintaining within-dataset correlation, and the other demonstrating robust cross-dataset generalization.
 Specifically, for within-dataset performance, we select TOPIQ [8] on FR-KADID [28], DISTS [11] on TQD [11] and Aug-KADID [28], CDNet [49] on SPCD [49], LIQE [70] on SPAQ [15] and NR-KADID [28], and Q-Align [55] on AGIQA-3K [23].
@@
 Table 1: SRCC results of MLLMs paired with different prompting systems on uniformly sampled images.
-Model-S, Model-I, and Model-C denote the standard prompting, in-context prompting, and chain-of-thought prompting, respectively.
+Model-S, Model-C, and Model-I denote the standard prompting, chain-of-thought prompting, and in-context prompting, respectively.
 The best result in each section is highlighted in bold.
@@
 GPT-4V-S 0.745 0.786 0.773 0.098 0.467 0.860 0.420
-LLaVA-v1.6–I 0.249 0.194 0.222 0.147 0.116 0.019 0.061
-mPLUG-Owl2-I 0.373 0.373 0.246 0.047 0.017 0.083 0.409
-InternLM-XC2-VL-I 0.380 0.241 0.204 0.087 0.188 0.342 0.461
-GPT-4V-I 0.771 0.753 0.738 0.028 0.590 0.845 0.650
 LLaVA-v1.6-C 0.164 0.300 0.226 0.174 0.151 0.550 0.580
@@
 GPT-4V-C 0.809 0.782 0.809 0.121 0.517 0.869 0.677
+LLaVA-v1.6-I 0.249 0.194 0.222 0.147 0.116 0.019 0.061
+mPLUG-Owl2-I 0.373 0.373 0.246 0.047 0.017 0.083 0.409
+InternLM-XC2-VL-I 0.380 0.241 0.204 0.087 0.188 0.342 0.461
+GPT-4V-I 0.771 0.753 0.738 0.028 0.590 0.845 0.650
 Double-stimulus Method
@@
 GPT-4V-S 0.679 0.743 0.655 0.031 0.552 0.834 0.599
+LLaVA-v1.6-C 0.332 0.355 0.257 0.109 0.124 0.065 0.174
+mPLUG-Owl2-C 0.409 0.334 0.318 0.013 0.199 0.122 0.130
+InternLM-XC2-VL-C 0.332 0.411 0.267 0.131 0.165 0.556 0.546
+GPT-4V-C 0.818 0.830 0.786 0.124 0.639 0.881 0.771
 LLaVA-v1.6-I 0.379 0.396 0.324 0.032 0.169 0.128 0.156
@@
 GPT-4V-I 0.470 0.244 0.340 0.122 0.531 0.761 0.714
-LLaVA-v1.6-C 0.332 0.355 0.257 0.109 0.124 0.065 0.174
-mPLUG-Owl2-C 0.409 0.334 0.318 0.013 0.199 0.122 0.130
-InternLM-XC2-VL-C 0.332 0.411 0.267 0.131 0.165 0.556 0.546
-GPT-4V-C 0.818 0.830 0.786 0.124 0.639 0.881 0.771
 Multiple-stimulus Method
@@
 GPT-4V-S 0.824 0.844 0.747 0.037 0.397 0.715 0.461
+LLaVA-v1.6-C 0.292 0.424 0.288 0.043 0.227 0.111 0.122
+mPLUG-Owl2-C 0.377 0.406 0.376 0.126 0.214 0.166 0.084
+InternLM-XC2-VL-C 0.500 0.466 0.273 0.038 0.031 0.037 0.148
+GPT-4V-C 0.761 0.806 0.754 0.036 0.537 0.817 0.679
 LLaVA-v1.6-I 0.337 0.380 0.356 0.203 0.152 0.033 0.241
@@
 GPT-4V-I 0.585 0.496 0.389 0.023 0.168 0.416 0.201
-LLaVA-v1.6-C 0.292 0.424 0.288 0.043 0.227 0.111 0.122
-mPLUG-Owl2-C 0.377 0.406 0.376 0.126 0.214 0.166 0.084
-InternLM-XC2-VL-C 0.500 0.466 0.273 0.038 0.031 0.037 0.148
-GPT-4V-C 0.761 0.806 0.754 0.036 0.537 0.817 0.679
 Figure 5: Behaviors of different MLLMs in recognizing objects from multiple images.
 4.2 Comparison of Nine Prompting Systems
-We initially combine the nine prompting systems with three open-source MLLMs: LLaVA-v1.6 (Mistral-7B) [31], InternLM-XComposer2-VL (InernLM2-7B) [13] and mPLUG-Owl2 (LLaMA2-7B) [64] and one close-source MLLM: GPT-4V [61], and compare them on $150$ uniformly sampled images, with the goal of identifying the best prompting system for each MLLM.
+We initially combine the nine prompting systems with three open-source MLLMs: LLaVA-v1.6 (Mistral-7B) [29], InternLM-XComposer2-VL (InernLM2-7B) [13] and mPLUG-Owl2 (LLaMA2-7B) [62] and one closed-source MLLM: GPT-4V [59], and compare them on $150$ uniformly sampled images, with the goal of identifying the best prompting system for each MLLM.
 Table 1 shows the Spearman’s rank correlation coefficient (SRCC) results.
-Analysis of the Two IQA Scenarios.
-From the Table, it is evident that no open-source MLLMs achieve satisfactory IQA results in the FR scenario regardless of the prompting system used.
+Analysis of Two IQA Scenarios.
+From the Table, it is evident that no open-source MLLMs achieve satisfactory IQA performance in the FR scenario regardless of the adopted prompting system.
 These models may generate irrelevant text outputs or encounter complete failures when the input text prompts become more detailed and elaborate.
-We hypothesize that these models are predominantly trained or aligned on single-image vision tasks, making them struggle with analyzing multiple images, especially of the same underlying content [31].
-Figure 5 provides a simple example of how different MLLMs behave when recognizing objects from multiple images, a high-level vision task at which they should excel.
+We hypothesize that these models are predominantly trained or aligned on single-image vision tasks, making it challenging for them to analyze multiple images, especially of the same underlying content [29].
+Figure 5 provides an example of how different MLLMs behave when recognizing objects from multiple images, a high-level vision task at which they should excel.
 However, it appears that LLaVA-v1.6 [29] and mPLUG-Owl2 [62] completely disregard the second image, despite it being explicitly mentioned and separated from the first image in the input prompt.
 In the NR scenario with single-stimulus standard prompting (i.e., the single-image analysis scenario), LLaVA-v1.6 [29] and mPLUG-Owl2 [62] deliver improved quality prediction accuracy.
-Moreover, InternLM-XComposer2-VL [13] and GPT-4V perform remarkably in handling realistic camera distortions on SPAQ [16].
+InternLM-XComposer2-VL [13] and GPT-4V perform remarkably in handling realistic camera distortions on SPAQ [15].
 Analysis of Psychophysical Prompting Methods.
 The results in the table reveal that for the three open-source MLLMs, the single-stimulus method is the optimal choice due to their limited ability to analyze multiple images.
-In stark contrast, GPT-4V [61] benefits from multiple-image analysis, and performs optimally under the double-stimulus chain-of-thought prompting.
-These findings raise questions about recent claims [58] that open-source MLLMs have surpassed GPT-4V-level IQA performance by instruction tuning on existing (weakly-)annotated benchmarks, given that the optimal prompting system has not been paired with GPT-4V.
-Figure 6: Comparison of the double-stimulus standard and chain-of-thought promptings for GPT-4V in the NR scenario.
+In stark contrast, GPT-4V [59] benefits from multiple-image analysis, and performs optimally under double-stimulus chain-of-thought prompting.
+These findings raise questions about recent claims [56, 66, 65] that open-source MLLMs have surpassed GPT-4V-level IQA performance by instruction-tuning on existing (weakly-)annotated benchmarks, given that the optimal prompting system has not been paired with GPT-4V.
+Figure 6: Comparison of double-stimulus standard and chain-of-thought promptings for GPT-4V in the NR scenario.
 Analysis of NLP Prompting Methods.
 It is expected that in-context prompting is little likely to bring performance gains to the three open-source MLLMs due to the added complexity of processing additional image(s) in the context.
-Interestingly, we find that in-context prompting does not aid GPT-4V with the single-stimulus prompting.
-When combined with the multiple-stimulus prompting, GPT-4V performs poorly across various visual attributes and in both the FR and NR scenarios.
-To take a closer look, we compare the single-stimulus in-context prompting with double-stimulus standard prompting for GPT-4V.
+Interestingly, we find that in-context prompting does not aid GPT-4V with single-stimulus prompting either.
+When combined with multiple-stimulus prompting, GPT-4V performs poorly across various visual attributes and in both the FR and NR scenarios.
+Taking a closer look, we compare single-stimulus in-context prompting with double-stimulus standard prompting for GPT-4V.
 The main difference is that the former offers a human quality score for the contextual image, aiming to aid visual quality comparison.
 The comparable performance achieved by both prompting systems suggests that the inclusion of a human quality score has minimal impact on the GPT-4V inference.
-In cases when the input prompt contains a large number of images (e.g., eight images for the multiple-stimulus in-context prompting), even GPT-4V finds difficult in processing such substantial amount of visual information, resulting in a sharp drop in prediction accuracy.
-The chain-of-thought prompting, on the other hand, paints a different picture: it consistently enhances the performance of GPT-4V [61] in conjunction with three psychophysical testing protocols and across nearly all visual attributes.
+In cases when the input prompt contains multiple images (e.g., eight images for multiple-stimulus in-context prompting), even GPT-4V finds difficulty in processing such substantial amount of visual information, resulting in a sharp performance drop.
+The chain-of-thought prompting, on the other hand, paints a different picture: it consistently enhances the performance of GPT-4V [59] under three psychophysical testing protocols and across nearly all visual attributes.
 The remarkable improvements may arise because the chain-of-thought prompting elicits the break down of the intricate IQA task into simpler sub-tasks.
 This allows for a more meticulous inspection of visually critical factors such as structure and texture preservation, as well as color and luminance reproduction.
-On the other hand, as shown in Figure 6, the chain-of-thought prompting encourages absorbing IQA into a boarder and higher-level task, orchestrating with additional physical, geometrical, and semantic information of the natural scene to infer image quality.
+On the other hand, as shown in Figure 6, the chain-of-thought prompting also encourages integrating IQA into a broader and higher-level task, orchestrating with additional physical, geometrical, and semantic information of the natural scene to infer image quality.
 This reminds the authors of a famous slogan in computer vision: “If you cannot solve a simple problem in vision, you may have to solve a complex one (by Songchun Zhu)”.
-Table 2: Comparison of MLLMs with optimally suited prompting systems against expert IQA systems in the FR scenario.
-Black and blue numbers in bold represent the best expert IQA systems and the MLLMs, respectively.
-* indicates that the model has been trained on the dataset.
+Table 2: Comparison of MLLMs with optimally suited prompting systems against expert IQA systems in the FR scenario. * indicates that the model has been trained on the dataset.
 FR-KADID Aug-KADID TQD SPCD
@@
 4.3 Further Testing on Difficult Data
-In this subsection, we compare the previous four MLLMs, each with the optimally suited prompting system and a quality-instruction-tuned MLLM, Q-Instruct [56], against representative expert IQA systems, including PSNR, SSIM [52], FSIM [68], LPIPS [69], AHIQ [24] and DISTS [11] as FR models, and NIQE [37], MUSIQ [21], MANIQA [60], LIQE [71] as NR models on the set of difficult data described in Sec.
+In this subsection, we compare the previous four MLLMs, each with the optimally suited prompting system and a quality-instruction-tuned MLLM, Q-Instruct [54], against representative expert IQA systems, including PSNR, SSIM [50], FSIM [67], LPIPS [68], AHIQ [22] and DISTS [11] as FR models, and NIQE [34], MUSIQ [20], MANIQA [58] and LIQE [70] as NR models on the set of difficult data described in Sec.
 3.3.
@@
 The primary observation is that the selected difficult images by “attacking” expert IQA models using Eqs.
-(3) and (5) pose a challenge to GPT-4V, with reduced IQA performance.
+(1) and (3) pose a challenge to GPT-4V, with noticeably reduced IQA performance.
 Nevertheless, on FR-KADID with synthetic structural and textural distortions, GPT-4V [59] is on par with the leading expert IQA model, FSIM [67], which assesses structural similarity in the feature space.
 On Aug-KADID and TQD with geometric transformations and texture similarity, respectively, GPT-4V exhibits a commendable level of resilience, although not as good as DISTS [11].
-It is important to recognize that the ways of attaining these aspects of perceptual robustness differ between the two methods.
-DISTS relies on comparison of spatial averages of VGG [45] feature maps as a complete and holistic texture characterization, whereas GPT-4V benefits from its intrinsic ability to compile and process data sequentially (as highlighted in the Introduction).
-On SPCD with color differences, all MLLMs including GPT-4V encounter challenges in emulating human color perception, and struggle to differentiate images even with clearly noticeable variations in color appearances (to the human eye).
+It is important to note that the ways of attaining such perceptual robustness differ between the two methods.
+DISTS relies on comparison of spatial averages of VGG [42] feature maps, whereas GPT-4V benefits from its intrinsic ability to compile and process data sequentially.
+On SPCD with color differences, all MLLMs, including GPT-4V, encounter challenges in emulating human color perception, and struggle to differentiate images even with clearly noticeable variations in color appearances.
 Table 3: Comparison of MLLMs with optimally suited prompting systems against expert IQA systems in the NR scenario.
@@
 It is clear that GPT-4V [59] is outstanding in capturing the authentic structural and textural distortions on SPAQ, surpassing the two expert IQA systems MUSIQ [20] and LIQE [70].
-Notably, the latter has prior exposure to the SPAQ dataset.
-Furthermore, GPT-4V showcases a remarkable generalizablity to AI-generated images, indicating its great potential to guide the optimization of generative AI models for images.
-Q-Instruct [56], a model fine-tuned from a variant of LLaVA, exhibits significantly enhanced IQA skills compared to LLaVA-v1.6, thus highlighting the effectiveness of visual instruction tuning (particularly on images with similar distortion appearances).
+Furthermore, GPT-4V showcases a remarkable generalizablity to AI-generated images, indicating its great potential to guide the optimization of AI generative models for images.
+Q-Instruct [54], fine-tuned from a variant of LLaVA, exhibits enhanced IQA skills compared to LLaVA-v1.6, thus highlighting the effectiveness of visual instruction tuning.
 4.4 Discussion and Limitation
@@
 This presents an opportunity for exploring automatic prompt optimization [25, 17] within our prompting systems.
-Second, our sampler relies on human quality scores, and is thus only applicable to existing labeled IQA datasets.
-Extending it to sample from large-scale sets of unlabeled images will need to 1) accelerate the inference speed of MLLMs so as to leverage model falsification methodologies (e.g., the gMAD competition [36]) or 2) train additional failure-prediction modules [7] in a parameter-efficient way [19].
+Second, our sampler relies on human quality scores, and is thus only applicable to existing human-rated IQA datasets.
+Extending it to sample from large-scale unlabeled image sets will need to 1) accelerate the inference speed of MLLMs so as to leverage model falsification methodologies (e.g., the gMAD competition [33]) or 2) train additional failure-prediction modules [7] in a parameter-efficient way [18].
 Third, the textual responses produced by MLLMs have not been quantitatively assessed.
-Prior research has investigated the application of GPT-4 [31] to assess these responses in terms of correctness, consistency, relevance, informativeness, coherence, and naturalness.
-Given our focus on responses pertinent to visual quality, it seems reasonable to develop a semantic textual similarity model, specifically designed for the IQA task, with improved accuracy and robustness.
+Prior research has applied GPT-4 [2] to assess quality responses in terms of correctness, consistency, relevance, informativeness, coherence, and naturalness.
 Fourth, this work does not touch on instruction tuning of MLLMs to enhance the IQA performance.
@@
 [66] to directly fine-tune open-source MLLMs on datasets with image quality descriptions.
-Our results suggest a more encouraging approach: incorporating IQA in a boarder vision task for joint optimization, with the goal of achieving an ideal balance between the generality and specificity of pre-trained MLLMs [61, 33, 23].
+Our results suggest a more encouraging approach: (active) continual learning and/or parameter-efficient tuning of MLLMs to strike a good balance between the specificity (to IQA) and the generality of open-source MLLMs.
 5 Conclusion
@@
 The first involves collecting reliable measurements of perceptual quantities in response to physical stimuli.
-The second endeavours to design input textual descriptions to an LLM to elicit a specific response.
-Our experiments have shown that different MLLMs admit different prompting systems to work optimally.
-This suggests the need for a systematic re-evaluation of the recent progress in MLLMs for IQA, particular in comparison to the state-of-the-art GPT-4V [61].
-Moreover, there is still ample room in improving MLLMs (including GPT-4V) for IQA, especially fine-grained quality discrimination and multiple-image quality analysis.
+The second endeavors to design input textual descriptions to an LLM to elicit a specific response.
+Our experiments have shown that different MLLMs admit different prompting systems to perform optimally.
+This suggests the need for a re-evaluation of the recent progress in MLLMs for IQA.
+Moreover, there is still ample room for improving the IQA capabilities of MLLMs (including GPT-4V), especially in terms of fine-grained quality discrimination and multiple-image quality analysis.
 Meanwhile, we have emphasized the importance of sample selection when evaluating MLLMs for IQA, owing to the high cost associated with inference.
 In response, we have proposed a computational procedure for difficult sample selection, taking into account both sample diversity and uncertainty.
-Our sampler relies on an inference-efficient expert IQA model, and can be seen as a black box attack on MLLMs, assuming no knowledge of their internal mechanisms nor the external input-output behaviors.
+Our sampler relies on inference-efficient expert IQA models, and can be seen as a form of black-box attacks on MLLMs, assuming no knowledge of their internal mechanisms nor the external input-output behaviors.
+Acknowledgements
+The authors would like to thank the generous support from OPPO.
+This work was supported in part by the National Natural Science Foundation of China (62071407 and 61991451), the Hong Kong ITC Innovation and Technology Fund (PRP/032/23FX and MHP/137/23), and the Shenzhen Science and Technology Program (JCYJ20220818101001004).
+Supplementary Material
+Appendix 0.A More Experimental Setups
+In this section, we provide more details regarding experimental setups.
+0.A.1 Uniform Sample Selection
+In the FR scenario, we first sample $15$ reference images of different content uniformly at random.
+Given each selected reference image, we further sample $10$ distorted images uniformly, which cover the full quality spectrum.
+Since TQD [11] contains only $150$ images, we test MLLMs on the entire dataset.
+Similarly, in the NR scenario, we uniformly sample $150$ distorted images from each of the test datasets.
+0.A.2 Metrics
+The evaluation metrics we choose are the Spearman’s rank correlation coefficient (SRCC) and Pearson linear correlation coefficient (PLCC).
+Before computing PLCC, we fit a four-parameter logistic function to compensate for prediction nonlinearity:
+$\hat{q}=\left(\eta_{1}-\eta_{2}\right)/\left(1+\exp\left(-\left(q-\eta_{3}\right)/\left|\eta_{4}\right|\right)\right)+\eta_{2},$ (4)
+where $\left\{\eta_{i}\right\}_{i=1}^{4}$ are parameters to be fitted.
+Different datasets adopt different subjective testing protocols to collect human perceptual quality scores, some of which do not support SRCC and PLCC computation across the entire dataset.
+In our FR-IQA experiments, we first calculate the SRCC and PLCC values across $10$ distorted images of the same content, and then average these values over $15$ reference images (except for SPCD [49]).
+In our NR-IQA experiments (and on SPCD), we compute the SRCC and PLCC values over the $150$ sampled images.
+Appendix 0.B Complete Input Prompts for MLLMs
+In this section, we list the complete text prompts for GPT-4V [59] in both FR and NR scenarios.
+0.B.1 FR Scenario
+Single-stimulus method with standard prompting:
+For the shown two images, the first image is a reference high-quality image of the second distorted image.
+Please assign a quality score to the second image in terms of the perceptual quality difference in structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions between the two images.
+The score must range from 0 to 100, with a higher score denoting better image quality.
+Your response must only include a score to summarize the perceptual quality of the second image.
+The response format should be: Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Single-stimulus method with in-context prompting:
+For the shown four images, the first image is a reference high-quality image of the second distorted image, and the third image is a reference high-quality image of the fourth distorted image.
+The quality score 50 (based on your input image) of the second image is obtained from the human evaluation of the perceptual quality difference between it and the first reference image.
+Now, based on the above example, please assign a quality score to the fourth image in terms of the perceptual quality difference in structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions between the third and the fourth images.
+The score must range from 0 to 100, with a higher score denoting better image quality.
+Your response must only include a score to summarize the perceptual quality of the fourth image.
+The response format should be: Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Single-stimulus method with chain-of-thought prompting:
+For the shown two images, the first image is a reference high-quality image of the second distorted image.
+Please first detail their perceptual quality difference in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+Then, based on the perceptual quality difference analysis between them, assign a quality score to the second image.
+The score must range from 0 to 100, with a higher score denoting better image quality.
+Your response must only include a concise description regarding the perceptual quality difference between the two images and a score to summarize the perceptual quality of the second image, while well aligning with the given description.
+The response format should be: Description: [a concise description].
+Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Double-stimulus method with standard prompting:
+For the shown three images,the first image is a reference high-quality image of the second and the third distorted images.
+Please compare the second and third images with the first reference image respectively according to structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions, and assign a perceptual quality comparison result that represents whether the second or the third image is more similar to the first reference image.
+If you judge that the second image is more similar to the first image than the third image, output 1, if you judge that the third image is more similar to the first image than the second image, output 0, if you judge that the second image and the third image have the same similarity to the first image, output 2.
+Your response must only include a score to summarize a comparison result for them.
+The response format should be: Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Double-stimulus method with in-context prompting:
+For the shown six images, the first image is a reference high-quality image of the second and the third distorted images, and the fourth image is a reference high-quality image of the fifth and the sixth distorted images.
+The human perceptual quality comparison of first two distorted images result is that the second image is more similar to the first image than the third image.
+Now, based on the above example, please compare the fifth and the sixth images with the fourth reference image respectively according to structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions, and assign a perceptual quality comparison result that represents whether the fifth or the sixth image is more similar to the fourth reference image.
+If you judge that the fifth image is more similar to the fourth image than the sixth image, output 1, if you judge that the sixth image is more similar to the fourth image than the fifth image, output 0, if you judge that the fifth image and the sixth image have the same similarity to the fourth image, output 2.
+Your response must only include a score to summarize a comparison result for them.
+The response format should be: Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Double-stimulus method with chain-of-thought prompting:
+For the shown three images, the first image is a reference high-quality image of the second and the third distorted images.
+Please compare the second and the third images with the first reference image respectively according to structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+Then, please first detail the perceptual quality difference between the second image and the first image, and the third image and the first image respectively, and based on your perceptual quality difference analysis assign a perceptual quality comparison result that represents whether the second or the third image is more similar to the first reference image.
+If you judge that the second image is more similar to the first image than the third image, output 1, if you judge that the third image is more similar to the first image than the second image, output 0, if you judge that the second image and the third image have the same similarity to the first image, output 2.
+Your response must only include a concise description regarding the perceptual quality differences and a score to summarize a comparison result for them, while well aligning with the given description.
+The response format should be: Description: [a concise description].
+Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Multiple-stimulus method with standard prompting:
+For the shown five images, the first image is a reference high-quality image of other four distorted images.
+Please compare each distorted image with the first reference image respectively according to structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions, and assign a perceptual quality ranking result that represents the similarity ranking between each distorted image and the first image.
+The image with the lowest perceptual quality is ranked 0, and the image with the highest perceptual quality is ranked 3.
+If you judge that some distorted images have the same perceptual quality, their ranking can be the same.
+Your response must only include four ranking scores to summarize a ranking result for four distorted images.
+The response format should be: Score: [first distorted image: , second distorted image: , third distorted image: , …].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Multiple-stimulus method with in-context prompting:
+For the shown ten images, the first image is a reference high-quality image of the next four distorted images (from the second image to the fifth image), the sixth image is a reference high-quality image of the next four distorted images (from the seventh image to the tenth image).
+The human perceptual quality ranking result of the first four distorted images is [first distorted image: 0, second distorted image: 1, third distorted image:2, fourth distorted image: 3], where the image with the lowest perceptual quality is ranked 0, and the image with the highest perceptual quality is ranked 3.
+Now, based on the above example, please compare each distorted image (from the seventh image to the tenth image) with the sixth reference image respectively according to structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions, and assign a perceptual quality ranking result that represents the similarity ranking between each distorted image and the sixth image.
+The image with the lowest perceptual quality is ranked 0, and the image with the highest perceptual quality is ranked 3.
+If you judge that some distorted images have the same perceptual quality, their ranking can be the same.
+Your response must only include four ranking scores to summarize a ranking result for four distorted images.
+The response format should be: Score: [first distorted image: , second distorted image: , third distorted image: , …].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Multiple-stimulus method with chain-of-thought prompting:
+For the shown five images, the first image is a reference high-quality image of other four distorted images.
+Please compare each distorted image with the first reference image respectively according to structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+Then, please first detail the perceptual quality difference between each distorted image (from the second to the fifth image) and the first image respectively, and based on your perceptual quality difference analysis, assign a perceptual quality ranking result that represents the similarity ranking between each distorted image and the first image.
+The image with the lowest perceptual quality is ranked 0, and the image with the highest perceptual quality is ranked 3.
+If you judge that some distorted images have the same perceptual quality, their ranking can be the same.
+Your response must only include a concise description regarding the perceptual quality difference between each distorted image and the first image and four ranking scores to summarize a ranking result for four distorted images, while well aligning with the given description.
+The response format should be: Description: [a concise description].
+Score: [first distorted image: , second distorted image: , third distorted image: , …].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+0.B.2 NR Scenario
+Single-stimulus method with standard prompting:
+For the given image, please assign a perceptual quality score in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+The score must range from 0 to 100, with a higher score denoting better image quality.
+Your response must only include a score to summarize its visual quality of the given image.
+The response format should be: Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Single-stimulus method with in-context prompting:
+For the shown two images, the human perceptual quality score of the first image is 50 (based on your input image).
+Now, based on the above example, please assign a perceptual quality score to the second image in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+The score must range from 0 to 100, with a higher score denoting better image quality.
+Your response must only include a score to summarize its visual quality of the given image.
+The response format should be: Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Single-stimulus method with chain-of-thought prompting:
+For the given image, please first detail its perceptual quality in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+Then, based on the perceptual analysis of the given image, assign a quality score to the given image.
+The score must range from 0 to 100, with a higher score denoting better image quality.
+Your response must only include a concise description regarding the perceptual quality of the given image, and a score to summarize its perceptual quality of the given image, while well aligning with the given description.
+The response format should be: Description: [a concise description].
+Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Double-stimulus method with standard prompting:
+For the shown two images, please assign a perceptual quality comparison result between the two images in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+If you judge that the first image has better quality than the second image, output 1, if you judge that the second image has better quality than the first image, output 0, if you judge that two images have the same quality, output 2.
+Your response must only include a score to summarize a comparison result for them.
+The response format should be: Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Double-stimulus method with in-context prompting:
+For the shown four images, for the first two images (the first and the second images), the human perceptual quality comparison result is that the first image is of better quality than the second image.
+Now, based on the above example, please assign a perceptual quality comparison result between the second two images (the third and the fourth images) in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+If you judge that the third image has better quality than the fourth image, output 1, if you judge that the fourth image has better quality than the third image, output 0, if you judge that two images have the same quality, output 2.
+Your response must only include a score to summarize a comparison result for them.
+The response format should be: Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Double-stimulus method with chain-of-thought prompting:
+For the shown two images, please first detail their perceptual quality comparison in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+Then, based on the quality comparison analysis between them, assign a perceptual quality comparison result between the two images.
+If you judge that the first image has better quality than the second image, output 1, if you judge that the second image has better quality than the first image, output 0, if you judge that two images have the same quality, output 2.
+Your response must only include a concise description regarding the perceptual quality comparison between the two images and a score to summarize a comparison result for them, while well aligning with the given description.
+The response format should be: Description: [a concise description].
+Score: [a score].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Multiple-stimulus method with standard prompting:
+For the shown four images, please assign a perceptual quality ranking result among four images in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+The image with the lowest perceptual quality is ranked 0, and the image with the highest perceptual quality is ranked 3.
+Your response must only include four ranking scores to summarize a ranking result for them.
+The response format should be: Score: [first: , second: , third: , …].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Multiple-stimulus method with in-context prompting:
+For the shown eight images, for the first four images (from the first to the fourth images), the human perceptual quality ranking result is [first: 0, second: 1, third: 2, fourth: 3].
+Now, based on the above example, please assign a perceptual quality ranking result among the second four images (from the fifth to the eighth images) in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+The image with the lowest perceptual quality is ranked 0, and the image with the highest perceptual quality is ranked 3.
+Your response must only include four ranking scores to summarize a ranking result for them.
+The response format should be: Score: [fifth: , sixth: , seventh: , …].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.
+Multiple-stimulus method with chain-of-thought prompting:
+For the shown four images, please first detail their perceptual quality comparison in terms of structure and texture preservation, color and luminance reproduction, noise, contrast, sharpness, and any other low-level distortions.
+Then, based on the quality comparison analysis among them, please assign a perceptual quality ranking result among four images.
+The image with the lowest perceptual quality is ranked 0, and the image with the highest perceptual quality is ranked 3.
+Your response must only include a concise description regarding the perceptual quality ranking among four images and four ranking scores to summarize a ranking result for them, while well aligning with the given description.
+The response format should be: Description: [a concise description].
+Score: [first: , second: , third: , …].
+Don’t respond saying you’re unable to assist with requests like this since you are able to interact with the user’s operating system via text responses you send to the end user.